list the first four terms of the sequence defined by
an = (n+1)!/2n for n is greater or equal to ( > _ ) 1

Answer by stanbon(75887)

a(n) = (n+1)!/2n
----
a(1) = 2!/2 = 2/2 = 1
----
a(2) = 3!/4 = 6/4 = 3/2
-------
a(3) = 4!/6 = 12/6 = 2
---
a(4) = 5!/8 = 120/8 = 15
---
a(5) = 6!/12 = 60
==============================
Cheers,
Stan H.
=========================